OpenStudy (anonymous):

find the value of tan-1(1/sqrt3) express your answer in radians

OpenStudy (jagatuba):

\[-\frac{\tan(1)}{\sqrt(3)} radians\]

OpenStudy (anonymous):

your job is to find a number (angle ) in the interval \[[-\frac{\pi}{2},\frac{\pi}{2}]\] whose tangent is \[\frac{1}{\sqrt{3}}\]

OpenStudy (anonymous):

think about where sine and cosine have values that involve \[\sqrt{3}\] and make sure that \[\frac{\sin(x)}{\cos(x)}=\frac{1}{\sqrt{3}}\]

OpenStudy (anonymous):

could you guys also help me graph the function y=2cos(theta+pi/3) ?

OpenStudy (anonymous):

there are really only two places to look. it is either \[\frac{\pi}{6}\] or \[\frac{\pi}{3}\] because they are the only ones that involve \[\sqrt{3}\]

OpenStudy (anonymous):

looks just like cosine, but shifted left \[\frac{\pi}{3}\] units, and goes from -3 to 3 instead of from -1 to 1

OpenStudy (anonymous):

here is a nice picture http://www.wolframalpha.com/input/?i=3cos%28x%2Bpi%2F3%29

OpenStudy (jagatuba):

Graph y=2cos(theta+pi/3) http://www.wolframalpha.com/input/?i=graph+y%3D2cos%28theta%2Bpi%2F3%29